# Engineering morphogenesis of cell clusters with differentiable programming

### v1

@@ Abstract
Understanding the rules underlying organismal development is a major unsolved problem in biology. Each cell in a developing organism responds to signals in its local environment by dividing, excreting, consuming, or reorganizing, yet how these individual actions coordinate over a macroscopic number of cells to grow complex structures with exquisite functionality is unknown. Here we use recent advances in automatic differentiation to discover local interaction rules and genetic networks that yield emergent, systems-level characteristics in a model of development. We consider a growing tissue with cellular interactions are mediated by morphogen diffusion, differential cell adhesion and mechanical stress. Each cell has an internal genetic network that it uses to make decisions based on its local environment. We show that one can simultaneously learn parameters governing the cell interactions and the genetic network for complex developmental scenarios, including the symmetry breaking of an embryo from an initial cell, the creation of emergent chemical gradients,homogenization of growth via mechanical stress, programmed growth into a prespecified shape, and the ability to repair from damage. When combined with recent experimental advances measuring spatio-temporal dynamics and gene expression of cells in a growing tissue, the methodology outlined here offers a promising path to unravelling the cellular basis of development.
@@
Inverse design tasks are usually translated into optimization problems, wherein the objective is to minimize the discrepancy between a property of the simulation system state and its desired value. While sophisticated computational models of cell behavior exist [13, 14], their applicability to reverse engineering remains limited, particularly in high-dimensional parameter spaces where parameter sweeps and evolutionary searches become inefficient. Automatic differentiation (AD) algorithms, on the other hand, enable efficient calculation of sensitivities of observables with respect to input parameters. While originally formulated for training large neural networks [15, 16], the same infrastructure can be used to differentiate through simulations of physical processes, enabling gradient-based optimization as a viable alternative to tackle hard inverse problems of various scientific interest. Examples include the design of self-assembling materials [17], the establishment of data-driven discretizations for partial differential equations [18], nonequilibrium statistical mechanics [19] and the design of quantum devices [20]. Neural Cellular Automata [21] represented an initial demonstration of how these methods could be used to engineer collective behaviors inspired by biology in lattice models. Hiscock, instead, showed that AD could be used to successfully learn meaningful coupling in ODE models of genetic networks [22].
@@
In our model, a population of three cell types is subject to Brownian motion and interacts via a Morse pairwise potential. The well depth of the potential between a pair of cells is dictated by the concentration of homotypic and heterotypic “cadherins” in each cell. Each cell type produces its respective homotypic cadherin and a shared heterotypic cadherin, regulated by chemical signaling. To achieve a desired spatial organization, we constructed three loss functions based on pairwise distances between cells. For a lattice-like structure, we minimized the sum of pairwise distances of cells of different types (Fig 2a). For organizing the structure into three lobes, we minimized the sum of pairwise distances of cells of the same type (Fig 2b). To achieve a core-shell structure, we minimized the distance from each cell to a pre-specified radius based on the cell type (Fig 2c).
In each case, the optimization learned simple mechanisms of spatial control. The optimization process identified effective solutions by allowing cells to observe their own type and regulate the pairwise potential between cell types. For a lattice-like structure, cells upregulated their heterotypic cadherin and downregulated their homotypic cadherin. Conversely, to create a lobed structure, they upregulated their homotypic cadherin and downregulated their heterotypic cadherin. These patterns of homo- and heterotypic adhesion are consistent with observations from cell culture sorting experiments [31]. Achieving core-shell structures instead required a more complex regulatory network (Fig 2c).
@@
Figure 6: Chemical Regulation of Homeostasis. (a) An initially imbalanced state (in the middle) grows into an equally unbalanced final state with untrained network parameters (left) and into a balanced state with learned parameters (right) (b) Visualization of the learned gene network. (c) Visualization of the chemical concentrations at the final state of the optimized simulation. (d) Average over 100 realizations of final proportion of red cells at the end of simulations of increasing lengths. (e) Average over 100 realizations of final proportion of red cells at the end of simulations with increasing proportions of red cells in the starting state. In both plots the shaded area corresponds to average plus or minus one standard deviation.
@@ Discussion
Our results demonstrate the power of gradient-based optimization methods in discovering mechanisms within the space of physical, cell-based models for collective behavior. By taking gradients through a physical model of cells interacting written using differentiable code, we were able to learn decision-making programs within individual cells to orchestrate growth into target shapes or with target properties. Additionally, we were able to overcome challenges associated with differentiating through the discrete nature of cell division events and non-differentiable loss functions through reinforcement learning based approaches. The mechanisms that were learned took the form of gene regulatory networks, which readily offers a biological interpretation that could even inform experiments. The learned mechanisms were local in nature; cells did not have access to global positional information and had to instead infer it by sensing chemicals and mechanics and communicating with nearby cells through these physical channels.
Our results are a promising first step towards inferring mechanisms in existing multicellular systems or learning how to program them for desired behaviors. In particular, morphogenetic mechanisms frequently utilize local feedback to coordinate intricate forms, often through chemical and/or mechanical signals. Examples include a myriad of developmental systems, from the development of the wing imaginal disc in Drosophila to the periodic segmentation of blocks in somitogenesis. The development of robust and useful organoid systems also requires an understanding of spatiotemporal chemical and physical cues needed to guide the growth of stem cells. Physical models could be optimized to reproduce experimental data from time-lapse imaging of such developmental systems to gain insights into individual decision making functions that recapitulate complex behaviors. Since we have shown that gene regulatory networks can be programmed to drive complex self-organization, further incorporation of experimental knowledge or distillation of the learned networks could even be instantiated experimentally in synthetic circuits. Furthermore, the soft sphere model we have presented here is simply a choice; the physical details can be modified according to the biological context the model is designed to ask questions for. The optimization can also be adjusted accordingly depending on the question being asked: mechanical parameters could be optimized over to learn regulation of cellular potentials, or hyperparameters like the number of chemicals. Questions about minimal mechanisms, robustness to variation and perturbations, or growth principles could also be framed and investigated. Gene expression datasets could be incorporated to bridge the gap between known and unknown information in biological systems. Our work opens up a multitude of avenues for further explorations of the landscape of growth programs for desired emergent behaviors in living systems.

### v2

@@ Abstract
Understanding the rules underlying organismal development is a major unsolved problem in biology. Each cell in a developing organism responds to signals in its local environment by dividing, excreting, consuming, or reorganizing, yet how these individual actions coordinate over a macroscopic number of cells to grow complex structures with exquisite functionality is unknown. Here we use recent advances in automatic differentiation to discover local interaction rules and genetic networks that yield emergent, systems-level characteristics in a model of development. We consider a growing tissue with cellular interactions mediated by morphogen diffusion, differential cell adhesion and mechanical stress. Each cell has an internal genetic network that is used to make decisions based on the cell’s local environment. We show that one can simultaneously learn parameters governing the cell interactions and the genetic network for complex developmental scenarios, including the symmetry breaking of an embryo from an initial cell by creation of emergent chemical gradients, homogenization of growth via mechanical stress, programmed growth into a pre-specified shape, and the ability to repair from damage. When combined with recent experimental advances measuring spatio-temporal dynamics and gene expression of cells in a growing tissue, the methodology outlined here offers a promising path to unravelling the cellular bases of development.
@@
Inverse design tasks are usually translated into optimization problems, wherein the objective is to minimize the discrepancy between a property of the simulated system state and its desired value. While sophisticated computational models of cell behavior exist [13, 14], their applicability to reverse engineering remains limited, particularly in high-dimensional parameter spaces where parameter sweeps and evolutionary searches become inefficient. Automatic differentiation (AD) algorithms, on the other hand, enable efficient calculation of sensitivities of observables with respect to input parameters. While originally developed for training large neural networks [15, 16], the same technical infrastructure can be used to differentiate through simulations of physical processes, offering gradient-based optimization as a viable alternative to tackle hard inverse problems of various scientific interest. Successful examples include the design of self-assembling materials [17], the establishment of data-driven discretizations for partial differential equations [18], nonequilibrium statistical mechanics [19] and the design of quantum devices [20]. Neural Cellular Automata [21] represented an initial demonstration of how these methods could be used to engineer collective behaviors inspired by biology in lattice models. Hiscock, instead, showed that AD could be used to successfully learn meaningful couplings in ODE models of genetic networks [22].
@@
In our model, a population of three cell types is subject to Brownian motion and interacts via a Morse pairwise potential. The well depth of the potential between a pair of cells is dictated by the concentration of homotypic and heterotypic “cadherins” in each cell. Each cell type produces its respective homotypic cadherin and a shared heterotypic cadherin, regulated by chemical signaling. Each cell type utilizes the same gene network, but can only secrete chemicals and express cadherins of its own type. To achieve a desired spatial organization, we construct three loss functions based on pairwise distances between cells. For a lattice-like structure, we minimize the sum of pairwise distances between cells of different types (Fig 2a). To organize the structure into three lobes, we minimize the sum of pairwise distances between cells of the same type (Fig 2b). To achieve a core-shell structure, we minimize the distance of each cell to a pre-specified radius, based on the cell type (Fig 2c).
The optimization process identifies effective solutions by allowing cells to observe their own type and regulate the pairwise potential between cell types. For a lattice-like structure, cells up-regulate their heterotypic cadherin and down-regulate their homotypic cadherin. Conversely, to create a lobed structure, they up-regulate their homotypic cadherin and down-regulate their heterotypic cadherin. These patterns of homo- and heterotypic adhesion are consistent with observations from cell culture sorting experiments [31]. Achieving core-shell structures requires a more complex regulatory network (Fig 2c).
@@
Figure 6: Chemical Regulation of Homeostasis. (a) An initially imbalanced state (in the middle) grows into an equally unbalanced final state with untrained network parameters (left) and into a balanced state with learned parameters (right) (b) Visualization of the learned gene network. (c) Visualization of the chemical concentrations at the final state of the optimized simulation. (d) Average over 100 realizations of final proportion of red cells at the end of simulations of increasing lengths. (e) Average over 100 realizations of final proportion of red cells at the end of simulations with increasing proportions of red cells in the starting state. In both plots the shaded area corresponds to average plus or minus one standard deviation. (f) Visualization of learned gene networks when each cell type can learn its own network.
@@ Discussion
Our results demonstrate the power of gradient-based optimization methods in discovering mechanisms within the space of physical, cell-based models for collective behavior. By taking gradients through a physical model of cells interacting written using differentiable code, we were able to learn decision-making programs within individual cells to orchestrate growth into target shapes or with target properties. Additionally, we were able to overcome challenges associated with differentiating through the discrete nature of cell division events and non-differentiable loss functions through reinforcement learning based approaches. The mechanisms that were learned took the form of gene regulatory networks, which readily offer a biological interpretation. The learned mechanisms were local in nature; cells did not have access to global positional information and had to instead infer it by sensing chemicals and mechanics and communicating with nearby cells through these physical channels.
Although the gene networks optimized here are too large to instantiate experimentally, they could be distilled to find minimal networks or constrained to incorporate prior biological knowledge. Furthermore, the soft sphere model we have presented here is simply one of the possible choices. The physical details can be modified to optimize over an appropriate forward model with the same methods we demonstrated in this work. Questions about minimal mechanisms, robustness to variation and perturbations, number of unique solutions or growth principles could also be framed and investigated within the same setting.
Our results are a promising first step towards inferring mechanisms in existing multicellular systems or learning how to program them for desired behaviors. In particular, morphogenetic mechanisms frequently utilize local feedback to coordinate intricate forms, often through chemical or mechanical signals. Examples include a myriad of developmental systems, from the development of the wing imaginal disc in Drosophila to the periodic segmentation of blocks in somitogenesis. The development of robust and useful organoid systems also requires an understanding of spatiotemporal chemical and physical cues needed to guide the growth of stem cells. Physical models could be optimized to reproduce experimental data from time-lapse imaging of such developmental systems to gain insights into individual decision making functions that recapitulate complex behaviors. Gene expression datasets could be incorporated to bridge the gap between known and unknown information in biological systems. Our work opens up a multitude of avenues for further explorations of the landscape of growth programs for desired emergent behaviors in living systems.